# HSIDMamba: Exploring Bidirectional State-Space Models for Hyperspectral Denoising

### v1

@@ Abstract
Effectively discerning spatial-spectral dependencies in HSI denoising is crucial, but prevailing methods using convolution or transformers still face computational efficiency limitations. Recently, the emerging Selective State Space Model(Mamba) has risen with its nearly linear computational complexity in processing natural language sequences, which inspired us to explore its potential in handling long spectral sequences. In this paper, we propose HSIDMamba(HSDM), tailored to exploit the linear complexity for effectively capturing spatial-spectral dependencies in HSI denoising. In particular, HSDM comprises multiple Hyperspectral Continuous Scan Blocks, incorporating BCSM(Bidirectional Continuous Scanning Mechanism), scale residual, and spectral attention mechanisms to enhance the capture of long-range and local spatial-spectral information. BCSM strengthens spatial-spectral interactions by linking forward and backward scans and enhancing information from eight directions through SSM, significantly enhancing the perceptual capability of HSDM and improving denoising performance more effectively. Extensive evaluations against HSI denoising benchmarks validate the superior performance of HSDM, achieving state-of-the-art results in performance and surpassing the efficiency of the latest transformer architectures by $30\%$.
@@ I Introduction
Hyperspectral images(HSI) refers to a type of remote sensing data that captures surface features across a broader spectrum of wavelengths. HSI includes the spectral characteristics of objects reflecting or emitting light of different wavelengths, and its unique spectral characteristics enable it to store more comprehensive information. Therefore, HSI has applications in various fields, including remote sensing, object detection, classification, image processing, and more[1, 2, 3, 4, 5].
The presence of noise in HSI is a multi-faceted issue, stemming from an array of sources including sensor imperfections, atmospheric disturbances, extraneous light reflections from surfaces, and malfunctioning of the capturing equipment. These pervasive noise contributors significantly degrade the quality and precision of hyperspectral data, thereby undermining its dependability for real-world applications. As a result of such noise contamination, the integrity of data-driven outcomes, including classification, localization, and analytical tasks, is compromised. This can have detrimental consequences on the veracity and efficacy of decision-making processes that rely on HSI. In recognition of these challenges, the imperative to efficiently suppress noise from HSI is underscored.
In contrast to traditional RGB and grayscale images, HSI contains a wealth of spectral information, posing two significant challenges: substantial redundancy in the spectral domain and computational burden associated with capturing long-range contextual dependencies. Prior assumptions about hyperspectral noise constrain traditional model-based approaches and require adjustments of different hyperparameters for varying scenes[6, 7, 8]. These limitations constrain their performance on real-world datasets and necessitate significant computational and temporal resources. Conversely, recent efforts leveraging deep learning networks have demonstrated superiority in hyperspectral denoising tasks, notably through architectures designed around convolutional neural networks(CNN) and transformer frameworks.
@@
Figure 1: Quantitative comparisons (PSNR (dB) vs. Inference Cost(second) of our HSDM against other deep learning methods on ICVL datasets under mixture noise.
While the prospects for hyperspectral image restoration based on CNNs are promising, existing work in this domain also encounters several challenges. Firstly, the weight-sharing strategy inherent in CNNs constrains the model’s ability to dynamically adapt to different inputs, potentially affecting its performance when dealing with unseen low-quality images. Secondly, the limited receptive field of individual convolutional layers, constrained by the size of the kernel, may significantly increase local restoration biases. Although some studies have proposed different convolution designs, current CNN-based restoration methods are still constrained by the effective receptive field[9, 10].
Subsequent Transformer architectures effectively addressed the limitations of receptive fields, leading to significant advancements in long-range dependency modeling and model performance. However, the inherent computational and memory complexity of the self-attention mechanism employed by Transformers results in quadratic growth in computational burden and memory requirements as sequence length increases[11, 12, 13].
@@
State Space Models (SSMs) have attracted significant research interest in addressing the challenge of modeling long sequences. They can be viewed as a fusion of recurrent neural networks (RNNs) and CNNs. Notable contributions in this domain include the Structured State Space (S4)[14] and its various extensions[15, 16, 17]. The recent development, Mamba[18], further enhances the S4 framework by incorporating a selection mechanism, enabling the model to focus on relevant information in an input-dependent manner. Additionally, hardware-aware optimizations are applied to ensure efficient training and inference. Mamba surpasses Transformers in handling large-scale datasets and exhibits linear scalability with sequence length, positioning it as a promising alternative for language modeling tasks. Concurrent research efforts have extended this architecture beyond 1D language tasks to 2D visual domains[19, 20, 21], including tasks such as image classification and biomedical image segmentation. However, to our knowledge, there has been no exploration of leveraging this efficient architecture for addressing HSI denoising.
Given the inherent long sequences generated by HSI, which necessitate the capabilities of Mamba, we propose Hyperspectral Denoising Mamba(HSDM), an effective model based on the selective state space model. HSDM harmoniously eliminated the shortcomings of convolution and attention mechanisms, enabling better modeling of spatial-spectral long-range dependencies in linear complexity. Moreover, HSDM introduces a multi-bidirectional continuous scanning mechanism(BCSM) tailored for HSI. Specifically, BCSM efficiently integrates information from multiple directions processed by designing forward and backward contiguous flattening SSM mechanisms, enhancing the performance of HSDM through residual and spectral attention mechanisms. As illustrated in Figure 1, HSDM achieves state-of-the-art performance in denoising and inference speed, demonstrating a well-balanced approach.
Experimental outcomes across diverse datasets validate the superior performance of our proposed HSDM compared to existing methodologies, evidenced by both objective metrics on synthetic datasets and visual quality assessments on real datasets. To summarize, the key contributions of our study are outlined below:
@@
  • We are the first to introduce a simple and effective Bidirectional State-Space Model, thereby devising a novel architecture distinct from CNNs and transformers for HSI denoising.
  • Addressing the inherent spatial and spectral correlations in HSI data, we propose a novel Continuous Bidirectional Scanning Mechanism to explore tailored variations of the Mamba framework.
@@
  • Extensive experiments demonstrate that our proposed MambaHSID achieves state-of-the-art performance and is $30\%$ faster than the latest transformer method, offering a more balanced solution in terms of performance and efficiency for HSI denoising.
@@ II Related Work
II-A Convolutional Neural Networks
CNNs have attracted significant attention for their ability to automatically extract noise patterns without prior knowledge of noise types. [22] leverage 2D convolutions effectively for HSI denoising, showing strong performance in both spatial and spectral domains. Additionally, some approaches utilizing 3D convolutions, exemplified by the works[23, 9], excel in capturing spatial-spectral structural information in high-dimensional data. Among these methods, QRNN3D[10], stands out for effectively modeling spatial-spectral dependencies and global spectral correlations, fully exploiting the interplay between spatial and spectral dimensions. Additionally, [24] employs a non-local neural network with end-to-end model assistance, incorporating considerations of both spectral low-rankness and spatial depth priors to mitigate HSI noise. Nonetheless, despite their impressive noise reduction capabilities, CNNs inherently local receptive fields may restrict their ability to capture long-range dependencies within the data.
II-B Vision Transformer
The Vision Transformer (ViT) has gained traction in computer vision, employing a unique architecture from traditional CNNs by leveraging the Transformer model initially designed for natural language processing. ViT excels across various tasks such as image classification, object detection, and segmentation due to its ability to handle images of any size and exhibit strong generalization capabilities.
ViT divides input images into patches and transforms the pixel data of each patch into a one-dimensional sequence, which is processed through multiple Transformer layers. The self-attention mechanism enables complex dependencies between pixels, capturing global information within the image.
In recent research on hyperspectral imaging with the Transformer architecture, innovative approaches have been proposed. For instance, [25] introduces a grouped pixel embedding module for HSI classification within local spatial-spectral contexts. [26] customize a mask-guided mechanism for effective HSI super-resolution, while [27] enhance denoising capabilities using window attention modules and global spectral attention to focus on spatial similarity and spectral dependencies.
However, despite its aptitude for capturing global interactions, Transformer suffers computational and data efficiency hampered by its requirement for large-scale datasets for training and its tendency to be computationally intensive, especially when dealing with high-resolution HSI.
@@ II-C State Space Models
Recent developments in state space models (SSMs) have shown a marked proficiency in encapsulating the dynamic interconnections and dependencies prevalent in language sequences. The S4 model, which specifical engineered to adeptly navigate long-range dependencies without sacrificing linear computational complexity, represents a significant stride in this research area. Expanding on these gains, the Mamba framework [18] introduces an SSM layer that adapts to data and a parallel scanning selection mechanism, which collectively exhibits enhanced capability in processing extensive sequences with linear complexity—a considerable advantage over attention-based transformers as noted in [12, 13].
In the domain of visual processing, SSMs have equally made impactful inroads. Applications of SSMs have been applied to visual tasks such as image classification [28] as well as in discerning long-range temporal dependencies within movie clip classification scenarios [29]. The advent of methodologies has heralded a period of innovation for state space models. Notably, the Mamba model distinguishes itself with its cutting-edge selective SSM feature, attaining an equipoise of linear computational complexity with the capacity to model long-term dynamics effectively. Our proposed HSDM innovation enhances the original Mamba model by optimizing its data-dependent state space layer for advanced HSI denoising efficiency and effectiveness.
@@ III Methods
Figure 2: Overall of the continuous spatiotemporal scanning mechanism. Sequentially passing through forward and backward continuous scanning mechanisms not only saves space and training time for the model but also better explores the spatial-spectral continuity of HSI.
In this section, we first introduce the basic concepts of state space models, the implementation details of mamba, and the details of our proposed HSDM architecture.
@@ III-A Preliminaries
State space models is a statistical model used to describe the evolution of a system over time. SSM typically models input data and hidden states for input sequences using Ordinary Differential Equations (ODE). A linear SSM can be represented as:
@@
Subsequently, Eq.1 are usually merged using a discretization process, which can be represented as follows:
$=\exp(\Delta\textbf{A})$ A ) (2)
$={(\Delta\textbf{A})^{-1}}(exp(\Delta\textbf{A})-I)\cdot\Delta{\textbf{B}}$ B
@@
After discretization, the discrete version Eq.1 can be rewritten as:
@@
In neural sequence modeling architectures, this discretized stack allows for learning the parameters of each layer through gradient descent. Recently, Selective State Space Models (termed Mamba) [29] have significantly enhanced the flexibility of SSMs by relaxing the time-invariance constraints on SSM parameters while preserving computational efficiency. Multiple works have demonstrated Mamba’s advantage in modeling long-range dependencies.
Bidirectional Selective SSM. By employing highly efficient parallel scanning, Mamba mitigates the impact of repetitive sequential nature. However, for hyperspectral data, the original Mamba one-way perception mechanism is not sufficient for analyzing hyperspectral data. Considering the continuity of spatial features, MambaHSID designed a bidirectional continuous Mamba scanning block to better model the sequence bi-directional by designing 8 different scanning starting points and directions. In this study, we explored Mamba’s continuous scanning mechanism for hyperspectral data by continuously processing 1D sequences using forward-backward SSM, to maximize the inductive bias of similarity in multidimensional data.
@@ III-B1 Overall Architecture
As illustrated in Figure 3, our proposed architecture of HSIDMamba(HSDM) comprises a shallow feature extractor followed by multiple layers of Hyperspectral Continuous Scan Layers (HCSL). The HCSL integrates a 2D spatial continuous scanning mechanism with a channel attention component. Initially, the HSI input undergoes feature extraction via the shallow feature extractor. Subsequently, these shallow features are fed into the HCSL. Through scaling and skip connections, the input undergoes 1D convolution before being passed into the Mamba-scan module, where the resulting anisotropic vectors are combined.
@@
where $Y_{i}n$ represents the input noise HSI, $K_{1}$ is a 3×3 convolutional kernel, $\otimes$ denotes the convolutional operation, and $F_{s},F^{\prime}$ represents the shallow feature, intermediate feature. After multi-layer extraction, a Denoiser is used to suppress HSI noise, represented as $Y_{out}$.
Figure 3: Overall framework of our proposed Hyperspectral Denoising Mamba. Adopting the holistic architecture of ViT, we integrate bidirectional continuous scanning mechanism, BSSM (b), and spectral attention (c) through the core component, HCSB (a).”
III-B2 Hyperspectral Continuous Scan Block
Upon entering the HCSL, the shallow features remain in their 2D representation. The overall network algorithm flowchart is shown in Algorithm1(a). To further process the data, we apply layer normalization for rearrangement, unfolding it into 1D data. Subsequently, this 1D data is passed through a 1D convolutional layer before entering the BSSM (Bidirectional State Space Model) layer.
$F_{t}mp=BSSM(LN(F_{in}))+F_{in}$ (5)
$F_{out}=CA(Conv(LN(F_{t}mp)))+F_{tmp}$
As depicted in Figure 2, to conduct a comprehensive analysis of state selection outcomes, we aggregated the results obtained from various scans, employing two rounds of normalization. For an intermediate feature derived from the Hyperspectral Continuous Scan Block(HCSB), denoted as $F_{i}^{\prime}$, the subsequent output feature $F_{i+1}^{\prime}$ of the HCSB can be precisely articulated as follows:
$F_{j}=Fold(rearrange(F_{i}^{\prime})),j=0,1,...,t$ (6)

### v2

@@ Abstract
Effectively modeling global context information in hyperspectral image (HSI) denoising is crucial, but prevailing methods using convolution or transformers still face localized or computational efficiency limitations. Inspired by the emerging Selective State Space Model (Mamba) with nearly linear computational complexity and efficient long-term modeling, we present a novel HSI denoising network named HSIDMamba (HSDM). HSDM is tailored to exploit the capture of potential spatial-spectral dependencies effectively and efficiently for HSI denoising. In particular, HSDM comprises multiple Hyperspectral Continuous Scan Blocks (HCSB) to strengthen spatial-spectral interactions. HCSB links forward and backward scans and enhances information from eight directions through the State Space Model (SSM), strengthening the context representation learning of HSDM and improving denoising performance more effectively. In addition, to enhance the utilization of spectral information and mitigate the degradation problem caused by long-range scanning, spectral attention mechanism. Extensive evaluations against HSI denoising benchmarks validate the superior performance of HSDM, achieving state-of-the-art performance and surpassing the efficiency of the transformer method SERT by $31\%$.
@@ I Introduction
Hyperspectral images (HSI) capture highly discriminative surface features over an extensive spectrum and retain complex spectral information enabling its application in diverse tasks, such as remote sensing, object detection, and image analysis [1, 2, 3, 4, 5]. Nevertheless, the ubiquity of noise—originating from sensor defects, atmospheric interference, and equipment malfunction, compromises the reliability and fidelity of HSI, impairing applications reliant on accurate spectral data.
As shown in Figure 2, HSI noises manifest in various forms, such as Gaussian, stripe, deadline, impulse, speckle, Poisson noise, and a mixture of the above. As a result of such noise contamination, the integrity of data-driven outcomes, including classification, localization, and analytical tasks, is compromised. This can have detrimental consequences on the veracity and efficacy of decision-making processes relying on HSI. In recognition of these challenges, the imperative to efficiently suppress noise from HSI is underscored.
Hyperspectral images surpass RGB and grayscale imagery with spectral richness but grapple with substantial spectral redundancy and computational demand from contextual dependencies. Despite model-based methods [6, 7, 8] leading to a presumption of hyperspectral noise yield satisfactory results, they require constant fine-tuning of superparameters across different scenes and demand substantial computational and temporal resources, thus constraining their real-world implementation. Convolutional Neural Networks (CNN) based deep learning techniques [9, 10] have shown potent potential in HSI denoising, despite inherent challenges such as static adaptability and local restoration biases due to the limited receptive field constrained by kernel size. On the other hand, Transformer architectures effectively addressed the limitations of receptive fields, leading to significant advancements in long-range dependency modeling and model performance [11, 12, 13]. However, the inherent computational and memory complexity of the self-attention mechanism employed by Transformers results in quadratic growth in computational burden and memory requirements as sequence length increases.
@@
Figure 1: Quantitative comparisons (PSNR (dB) vs. Inference Cost(second)) of our HSDM against other deep learning methods on ICVL datasets under mixture noise.
Figure 2: Performance of different noises. The visual image is synthesized by bands 29, 19, and 9 of flowers_msflowers\_ms in CAVE.
@@
Recently, significant research interest has been vested in State Space Models (SSMs) due to their potential to confront the challenge of modeling long sequences. They can be viewed as a fusion of recurrent neural networks (RNNs) and CNNs. Notable contributions in this domain include the Structured State Space (S4) [14] and its various extensions [15, 16, 17]. The recent development, Mamba [18], not only enhances the S4 framework by incorporating a selection mechanism for an input-dependent focus on relevant information but also ensures efficient training and inference through hardware-aware optimizations. Mamba exhibits linear scalability with sequence length, positioning it as a promising alternative for language modeling tasks. Concurrent research efforts have extended this architecture beyond 1D language tasks to 2D visual domains, including tasks such as image classification and image segmentation [19, 20, 21]. Despite this, it is still challenging to directly adapt the original Vision Mamba [22] architecture to HSI denoising because the high dimension and abundant spectral data dramatically increase the parameters of the model and demand extensive computational resources. To the best of our knowledge, the potential of leveraging this efficient architecture for addressing HSI denoising challenges has not yet been explored.
To address these issues, we propose a network designed for the challenges posed by the inherently long sequences found in HSI. Our solution, termed the Hyperspectral Image Denoising Mamba-based network (HSDM), leverages a selective state space model. HSDM introduces an innovative Hyperspectral Continuous Scanning Block (HCSB) that is meticulously crafted to strengthen spatial-spectral interactions and combat the noise characteristic efficiently. In particular, HSDM effectively overcomes the inherent drawbacks of both convolutional and attention mechanisms. This is achieved by implementing an advanced 2D Selective Scan Module that facilitates a more robust and comprehensive modeling of spatial-spectral long-range dependencies while maintaining linear computational complexity.
Compared to General Vision Mamba, which typically employs a sweep scanning strategy, HSDM utilizes an alternating continuous scanning mechanism of forward and backward passes. This mechanism achieves comprehensive coverage across all scanning vectors, and in doing so, it enhances computational efficiency and diminishes parameter requirements. The bidirectional continuous scanning mechanism allows for a smooth transition of features before scanning, minimizing potential distortions. Moreover, to further augment the performance of the Mamba network in HSI denoising, we have incorporated the widely proven concept of residual learning. Additionally, we employ a simple yet effective spectral attention module to identify and emphasize pivotal spectral features. This enhancement has equipped the HSDM to deliver superior noise reduction while preserving the integrity of salient image characteristics, as illustrated in Figure 1.
To summarize, the main contributions of this study are summarized as follows:
@@
  • We are the first to introduce a simple yet effective Bidirectional State-Space Model for HSI denoising, devising a novel architecture distinct from CNNs and transformers.
  • To strengthen spatial-spectral interactions and context representation learning, we propose the Bidirectional Continuous Scanning Mechanism to improve detailed information denoising ability effectively and efficiently.
@@
  • Extensive experiments demonstrate that HSDM achieves state-of-the-art performance and is $31\%$ faster than the transformer method SERT.
@@ II Related Work
In this section, we chiefly explore the latest developments in research on data-driven deep learning techniques. While we refrain from delving into detailed expositions of model-driven approaches, they will be included for comparative analysis in the experimental evaluations.
II-A CNN-based Hyperspectral Denoising
CNNs effectively extract noise in HSI under various noise conditions, while simultaneously suppressing noise across spatial and spectral dimensions. [23] leverage 2D convolutions effectively for HSI denoising, showing strong performance in both spatial and spectral domains. Additionally, 3D convolution significantly enhances the utilization of spatial-spectral structural information. Demonstrated by several techniques, notably in the studies by[24, 9], 3D convolutions have proven exceptionally adept at capturing the intricacies of spatial-spectral structural information within high-dimensional data. Among the discussed methods, TRQ3DNet[25] particularly excels in modeling the spatial-spectral dependencies and capturing global spectral correlations, thereby fully leveraging the synergy between spatial and spectral dimensions. Specifically, QRNN3D[10] achieves impressive results by combining 3D convolution and recurrent pooling functions to more effectively incorporate spatial-spectral correlations and global contextual information. Nonetheless, despite their formidable noise reduction capabilities, the inherently local receptive fields of CNNs may impede their ability to grasp long-range dependencies in HSI.
II-B ViT-based Hyperspectral Denoising
Vision Transformer (ViT) has introduced a paradigm shift in the realm of computer vision. Diverging from the traditional confines of CNN architectures, ViT capitalizes on the robust Transformer model, originally conceived for tackling natural language processing challenges. Significantly, the effectiveness of the Transformer’s attention mechanism for long-range modeling has been empirically validated in previous works, underscoring its potential utility in a variety of applications. Outfitted with the capacity to process images of varied sizes and demonstrate powerful generalization, ViT surpasses traditional methods in numerous tasks such as image classification, object detection, and segmentation. This paradigm shift mitigates the previously highlighted shortcomings of CNNs.
In recent research into HSI denoising, a variety of innovative methods utilizing the Transformer architecture have been advanced. [12, 13] introduce enhancements to the foundational attention mechanism of the Transformer model, incorporating design elements specifically tailored to the spectral module. These modifications not only accelerate processing speed relative to the original Transformer but also address spectral distortion and edge blurriness, which are common in long-range modeling of HSI data. [26] differentiates itself by mitigating these issues through the preservation of local sensitivity bias, thereby establishing a model that correlates image content for effective HSI denoising. Nevertheless, in spite of determined endeavors to augment performance, the Transformer model encounters constraints in computational and data efficiency due to its dependence on large-scale datasets for effective training, coupled with a computational burden that intensifies when processing high-resolution HSI.
Figure 3: Illustration of the 2D-Selective-Scan of General Vision Mamba and HSDM. Compared to generic scanning, our continuous scan strategy achieves smoother transitions by alternatingly merging forward and backward features into the final 2D feature map.
@@ II-C State Space Models
While CNNs and Transformer models have greatly advanced the field of deep learning through their respective abilities to handle spatial hierarchies and long-range dependencies, researchers continually seek improvements in both architectural efficiency and capability for sequence modeling. It is in this quest for advancement that recent developments in state space models (SSMs) have garnered attention. These models have demonstrated a remarkable proficiency in encapsulating dynamic interconnections and dependencies that are prevalent in language sequences, offering an alternative approach that complements the strengths. S4 [14], specifically designed to adeptly manage long-range dependencies while maintaining linear computational complexity, marks a considerable advancement in this area of research. Building upon these improvements, Mamba [18] introduces an innovative SSM layer capable of dynamically adapting to data. Additionally, it employs a parallel scanning selection mechanism, effectively addressing some of the well-known limitations associated with RNNs and SSMs. Collectively, these developments enable the processing of extensive sequences with linear complexity. This represents a significant advantage over attention-based transformers, particularly in scenarios where managing long-range dependencies is critical.
In the domain of visual processing, SSMs have equally made impactful inroads. Applications of SSMs have been applied to visual tasks such as image classification [27] as well as in discerning long-range temporal dependencies within movie clip classification scenarios [28]. The advent of methodologies has heralded a period of innovation for state space models. Notably, the Mamba model distinguishes itself with its cutting-edge selective SSM feature, attaining an equipoise of linear computational complexity with the capacity to model long-term dynamics effectively. Our proposed HSDM innovation enhances the original Mamba model by optimizing its data-dependent state space mechanisms for advanced HSI denoising efficiency and effectiveness.
@@ III Methods
In this section, we commence with an introduction to the foundational concepts underpinning state space models. Following this, we delve into an in-depth exploration of the implementation specifics of our proposed HSDM. Finally, we present a succinct analysis and discussion of the computational complexity inherent in various mainstream architectures.
@@ III-A Preliminaries
State space models are statistical models designed to quantify the temporal evolution of a system. In the realm of sequence data, SSMs usually model input data and corresponding hidden states using Ordinary Differential Equations (ODEs). With this unique capability, SSMs have become a pivotal tool in handling temporal data. To further clarify this concept, consider the following example of a linear SSM which can be represented mathematically as follows:
@@
Following the basic formula (Eq.1), a discretization process is often employed to integrate these equations. This process can be mathematically represented as:
$\overline{\textbf{A}}=\exp(\Delta\textbf{A})$ (2)
$\overline{\textbf{B}}={(\Delta\textbf{A})^{-1}}(exp(\Delta\textbf{A})-I)\cdot\Delta{\textbf{B}}$
@@
Applying a discretized structure in the neural sequence modeling architectures facilitates the intricate process of learning layer-specific parameters using gradient descent techniques. Upon applying the discretization process, the continuous equation (Eq.1), mute to the linearity of the system and finite changes in time, manifests into its discrete counterpart, articulated as:
@@
Mamba has markedly expanded the adaptability of SSM frameworks by attenuating the rigid constraints of time-invariance on the SSM parameters, simultaneously maintaining a high level of computational efficacy. This revolutionary approach has been underscored by numerous studies, illustrating Mamba’s proficiency in capturing and modeling long-range dependencies with remarkable accuracy, thereby setting a new benchmark for the effectiveness of SSMs.
@@ III-B1 Overall Architecture
The architecture of our proposed HSIDMamba (HSDM) model is depicted in Figure 4, and the overall network algorithm flowchart is shown in Algorithm1. Starting with an initial shallow feature extractor that teases out preliminary spectral-spatial characteristics from the input HSI. This is followed by an ensemble of Hyperspectral Continuous Scan Layers (HCSL) that form the core of the network. Each HCSL is engineered with a dual-faceted focus: a spatially-aware continuous scanning mechanism to capture the two-dimensional intricacies and a channel attention component that fine-tunes the spectral features for enhanced discriminative learning.
@@
where $Y_{i}n$ represents the input noise HSI, $K_{1}$ is a 3×3 convolutional kernel, $\otimes$ denotes the convolutional operation, and $F_{s},F^{\prime}$ represents the shallow feature, intermediate feature, respectively. After multi-layer extraction, a Denoiser is used to suppress HSI noise, represented as $Y_{out}$.
Once the HSI input is subjected to the shallow feature extraction phase, the emergent feature maps are channeled into the HCSLs. Within this ambit, a dynamic interaction of scaling and skip-connection strategies facilitates a transformative 1D convolution process. This process is pivotal in conditioning the features for the subsequent Mamba-scan module, which adeptly synthesizes the anisotropic vectors into a coherent representation, conducive to subsequent modeling tasks.
Figure 4: Overall framework of our proposed Hyperspectral Denoising Mamba. Adopting the holistic architecture of ViT, we integrate a bidirectional continuous scanning mechanism. (a) Hyperspectral Continuous Scan Block, alternatingly integrating features from forward and backward scanning directions. (b) BSSM, an augmented State Space module that expands the scanning directions, and (c) spectral attention, to enhance the utilization of spectral information.
III-B2 Bidirectional Continuous Scanning Module
Incorporating highly efficient parallel scanning, Mamba effectively reduces the limitations of a repetitive sequential nature. However, when addressing hyperspectral data, the original Mamba’s unidirectional perception and scanning mechanisms prove inadequate. Recognizing the need to capture the continuity of spatial features, HSDM introduces a Bidirectional Continuous Scanning Module (BCSM). The strategy enhances the scanning process by alternatingly integrating features from the forward and backward scanning directions.
$FS_{i}=\sum_{i=1}^{4}Scan(Direction_{i},StartPoint_{i})$ (5)
where $FS_{i}$ represents the forward scan from the $i^{th}$ starting point facilitated by its corresponding scanning direction. For the backward scan, the approach involves inverting the scanning direction relative to each original starting point, thus allowing for a comprehensive bidirectional modeling of the data sequence:
$BS_{i}=FS_{i}(Reverse(Direction_{i}))$ (6)
By integrating this bidirectional scanning methodology with the linearized sequence data, the approach not only significantly enhances the inductive bias for simulating the similarities across multidimensional data but also notably curtails both inference duration and memory demands. The entire scanning method is illustrated in Figure 3, providing a visual representation of the process.
III-B3 Hyperspectral Continuous Scan Block
HCSL consists of multiple alternately Hyperspectral Continuous Scan Blocks (HCSBs), each with dual scaling to selectively retain feature information. Inside the HCSL, the preserved features still hold their 2D form. Data undergo normalization and reordering layers for refinement, transforming features into a one-dimensional format. Before entering the Bi-directional State Space Model (BSSM) layer and the spectral attention mechanism, these one-dimensional data go under layer normalization for consistent scaling and better spectral dependency capture. The process can be expressed as:
$F_{tmp}=BSSM(LN(F_{in}))+Scale(F_{in})$ (7)
$F_{out}=CA(Conv(LN(F_{t}mp)))+Scale(F_{tmp})$
where scale operation here is realized by 3*3 convolution.
As depicted in Figure 3, we aggregated the results obtained from various scans, employing two rounds of normalization. For an intermediate feature derived from HCSB, denoted as $F_{i}^{\prime}$, the subsequent output feature $F_{i+1}^{\prime}$ of the HCSB can be precisely articulated as follows:
$F_{j}=SiLU(DWConv(Fold(rearrange(F_{i}^{\prime}))))$ (8)